What is the boiling point of a 0.66 m solution of a nonvolatile solute in benzene?

Constants: Kb = 2.53 oCm-1 Tb= 80.1 oC
Answer #1

Solution:

The elevation of boiling point (ΔT) is related with molality of solute (m) as,

ΔT = kb x m = 2.53 °C m-1 x 0.66 m

= 1.6698 °C = 1.7 °C

Boiling point of solution (Ts) is calculated from ΔT and Tb as,

ΔT = Ts - Tb

Ts = Tb + ΔT = 80.1 °C + 1.7 °C = 81.8 °C